differentiate the following function with respect to x
cos^-1(sqrt(1+x / 2),-1

Grade:12

1 Answers

Rinkoo Gupta
9 years ago

y=cos-1v(1-x2)



dy/dx=-1/v1-(1-x2) .{-2x/v(1-x2)}



=1/v(1-x2)

Thanks & Regards

Rinkoo Gupta

AskIITians Faculty